### v1

# Free Energy Calculations using Smooth Basin Classification

Abstract
The efficiency of atomic simulations of materials and molecules can rapidly deteriorate when large free energy barriers exist between local minima. We propose smooth basin classification, a universal method to define reaction coordinates based on the internal feature representation of a graph neural network. We achieve high data efficiency by exploiting their built-in symmetry and adopting a transfer learning strategy. We benchmark our approach on challenging chemical and physical transformations, and show that it matches and even outperforms reaction coordinates defined based on human intuition.
Dynamic simulations of materials and molecules can provide both qualitative and quantitative information on chemical and physical transformations. While explicit integration of Newton’s equation of motion requires timesteps in the order of a femtosecond, the time required for many transformations to complete is often beyond the micro- or millisecond range because of large energy barriers between initial and final state. As such, simulation of a single transformation can easily require billions of time steps [1, 2].
Over the past decades, a variety of advanced sampling techniques have been proposed in order to improve the sampling efficiency of dynamic simulations. A particularly intuitive approach is to bias the dynamics along certain directions in phase space, effectively lifting the system over obstructive free energy barriers. The bias energy is defined along a certain predefined reaction coordinate or collective variable (CV). The CV is a function of the atomic coordinates and should naturally discriminate between initial and final state without violating the physical symmetries in the system. Specifically, the CV should remain invariant with respect to global translations and rotations, as well as permutations of atoms which are chemically identical. In the vast majority of practical applications, such CVs are manually defined based on chemical and physical intuition, and they usually contain simple invariants such as specific unit cell parameters or coordination numbers of one chemical species with respect to another. However, manual definition can be highly challenging and time-consuming, because the metastable intermediate(s) and/or transition state(s) are not always known in advance, or because the required invariances are difficult to satisfy [3, 4, 5].
Considering the numerous achievements of machine learning in atomistic modeling, it is sensible to anticipate that suitable reaction coordinates can also be learned from (unbiased) simulation data, and ideally without requiring human intuition about the transition mechanism(s). A wide variety of such methods have been proposed over the past decade, and we briefly mention a few notable examples while referring the reader to specialized reviews for a more comprehensive discussion [6, 7]. Supervised methods learn a continuous discriminant function which separates free energy minima based on user-defined structural descriptors, for example using (deep) linear discriminant analysis [8, 9, 10]. Unsupervised methods do not directly learn a discriminant function but instead identify so-called slow modes, either within a given set of descriptors or entirely end-to-end; see e.g. time-lagged independent component analysis or time-lagged autoencoders [11, 12, 13]. While these methods demonstrate the potential of using machine learning for defining reaction coordinates, their widespread adoption has remained rather limited. We hypothesize that this is due to inherent limitations regarding the required volume of simulation data, the difficulty of constructing an appropriate set of input features, or the difficulty in interpreting and validating the obtained CVs.
In this work, we propose smooth basin classification (SBC); a universal method to construct CVs and compute free energy differences for chemical and physical transformations. SBC builds upon the successful development of graph neural networks (GNNs) as effective interatomic potentials by using their learned feature space as ansatz for constructing physically meaningful CVs [14, 15, 16]. We apply SBC for complex rare events in both materials science and chemistry, and demonstrate that it matches or outperforms CVs based on human intuition. Relative free energy predictions based on learned CVs are within 1 kJ/mol of reference values as computed based on established CVs for each of the case studies. In addition, we observe that SBC requires orders of magnitude less training data as compared to existing approaches [6, 7, 8, 9, 10, 11, 12, 13]. This is a direct consequence of the fact that the constructed CVs are embedded in the learned feature space of the GNN, thereby inheriting all of their established benefits regarding data efficiency and universality.
Consider a system with two basins of attraction, named A and B. Each point in the system’s coordinate space can be uniquely labeled with A or B depending on the trajectory of a geometry optimization; belongs to basin A whenever it optimizes to any of the local minima within A, and vice-versa for B (Figure 1, a). In SBC, we aim to construct a mapping between each configuration and its associated basin of attraction. Instead of defining this mapping directly in terms of the atomic coordinates – which are not invariant with respect to the symmetry operations in the system – we embed it into the learned feature space of a GNN which has been pretrained to reproduce the potential energy surface of the system (Figure 1, b). In general, the interaction layers in the GNN parse the chemical environment of atom into a feature vector in this space (which typically contains anywhere between 8 and 64 dimensions). Importantly, this feature vector is rigorously invariant with respect to all symmetry operations in the system, and is specifically trained to capture the physical interactions in the system. Both the atomic energy and the total energy are computed from the atomic feature vectors using a simple readout function:
Because the atomic features are maximally informative for determining the interaction energy of a given atomic environment, we may assume that they are also highly informative in predicting the basin of attraction, i.e. the final destination of an energy minimization. In full analogy with equation 1, we can therefore construct a prediction for the basin of attraction using the same GNN but with a different readout function (Figure 1, c):
in which we introduce the per-atom log probabilities as an invariant tuple which measures how strongly the model associates a given atomic environment to basin A or B – as commonly done in a classification setting [17]:
For example, if the model believes that a given atomic environment is more likely to appear in phase as opposed to phase , it will predict log probabilities such that . After summing the per-atom log probabilities to obtain the total log probabilities , the predicted basin of attraction of the total configuration is then determined by the phase which received the largest log probability in :
A more detailed overview of the model architecture can be found in the Supplemental Material (Section S1).
While the prediction of the basin of attraction is essentially a discrete classification task, the actual log probabilities are smooth functions of the atomic feature vectors and hence the atomic coordinates – which means that they can be used to introduce biasing forces and/or virial stress. Ideally, we would like to design this bias such that it can drive transitions from and to any of the basins. To this end, consider the difference between two log probabilities:
In essence, measures how much more likely the model thinks a given configuration belongs to basin B rather than basin A. It is therefore strictly positive in basin B and strictly negative in basin A provided that the phase readout is properly trained (Figure 1, d). Clearly, the quantity correctly discriminates between two free energy minima while also being continuously differentiable with respect to the atomic coordinates (using standard backpropagation). It is therefore, by definition, a valid collective variable for the transition between basins A and B. Note that the invariance of with respect to global translations and rotations as well as permutation of chemically identical atoms is guaranteed by construction of the GNN. The built-in symmetry of SBC towards translations, rotations, and permutations incorporates important physical behavior into the model, and should be contrasted with both common practice CVs and most other ML-based approaches [9, 13, 12, 22]. which often use functions and/or input features that depend explicitly on the atomic indices (thereby violating permutation invariance) or specific components of the unit cell vectors (thereby violating rotational as well as periodic invariance).
The phase readout layers are analogous to the existing energy readouts found in conventional GNNs [14, 15, 16] (Figure 1, c), and essentially consist of a multilayer perceptron (MLP) containing at most two nonlinearities and about 1000 weights in total (compared to 100,000 weights in the GNN). The classifier is trained while keeping the GNN and energy readouts fixed, similar to conventional transfer learning approaches. Its training data consists of atomic geometries which are labeled with their corresponding basin of attraction (e.g. A or B). To ensure that the classifier remains accurate for out-of-equilibrium geometries, we employ a combination of enhanced sampling (in ) and active learning to explore transition regions between phases in an iterative fashion (Figure 1, e). More information regarding model training and the active learning approach can be found in the Supplemental Material (Sections S1 and S2).
Figure 2 (a) shows a nonreactive phase transition in a solid-state system named MIL-53(Al), a flexible metal-organic framework based on a winerack topology [23]. The system exhibits two basins of attraction which are relevant for tuning the material towards gas storage and/or sensing applications; a closed pore () and open pore () phase. In these types of transitions, i.e. those for which the unit cell parameters vary significantly between the different basins, researchers often employ some combination of unit cell parameters as collective variable – in this case, the unit cell volume [24]. Figure 2 (a) demonstrates the performance of the unit cell volume as formal discriminator between the two basins, and compares it with a difference of learned log probabilities as proposed in equation 5. Both criteria perform excellent in terms of classification accuracy, and in addition exhibit a very strong correlation. For both collective variables, the misclassified geometries are all localized near the transition region (such they will not cause substantial differences in the relative stability). Note that the network learns to distinguish low- and high-volume configurations based on atomic environments only, i.e. without explicitly relying on unit cell parameters as input.
Using the log probability difference as CV (see equation 5), we can employ conventional enhanced sampling techniques and evaluate the free energy profile . Here, we choose to employ umbrella sampling with harmonic restraints on the CV:
in which we emphasize that the bias energy contains an implicit dependence on both the atomic coordinates and the unit cell parameters through the learned atomic environment features [25]. We use the multistate Bennett acceptance ratio (MBAR) equations to combine samples from multiple umbrellas in a statistically optimal manner and obtain free energy profiles as shown in Figure 2 (b, top) [18, 19, 20, 21]. The reference free energy profile using the unit cell volume as bias coordinate is added for completeness in Figure 2 (b, bottom). The total relative free energy difference between the and basin, as obtained by integrating each of the profiles in their respective minima, agrees to within 1 kJ/mol between both CVs. Note that while relative stabilities are intrinsic to the system, the actual free energy profiles depend on the particular choice of CV, so the learned and reference free energy profiles need not have the same shape (although there exist ways to transform one profile into another [26]).
For global phase transformations in solids such as the one in Figure 2, a size-extensive definition of as a sum over per-atom log probabilities is intuitively sensible because it gives the model the ability to assign nonzero log probability contributions to all atoms simultaneously. However, in chemical transformations, the rare event is most often a local process whose free energy barrier does not necessarily depend on the size of the surrounding environment (e.g. the catalyst). To demonstrate the applicability of our approach to those cases, we investigate a proton transfer reaction from a Brønsted acid site in a zeolite to an alkene guest (Figure 3, a). The reference collective variable in this case is a coordination number difference which is defined explicitly in terms of the indices of the oxygen atoms of the active site as well as the indices of the carbon atoms of the molecule and all of the hydrogens – see the Supplemental Material (Section S5) for more details. While such definitions are extremely common, they violate permutation symmetry and make each collective variable highly system-specific. Figure 3 (a) shows the classification performance of the coordination number versus a learned CV based on equation 5. Again, classification accuracies are excellent for both the proposed CV based on chemical intuition as well as the learned log probability difference. Misclassified geometries are localized at or near the transition region, which means that they will not contribute to the relative free energy of both phases. Indeed, as indicated in Figure 3 (b), the free energy difference between both basins as computed from either of the two profiles agrees to within 1 kJ/mol.
A final case study concerns the conformational free energy landscape of alanine dipeptide, given that it is the most commonly studied system in the context of free energy calculations and the automated development of collective variables [12, 27, 13, 28, 29, 30, 9, 31, 32]. At the PBE-D3(BJ) level of theory and room temperature conditions, the molecule is observed to exhibit two main stable conformations; the c7eq and c7ax conformations. There exist multiple transition paths between both conformations, and each comprises multiple potential energy minima. The free energy of the system is most often visualized using a two-dimensional free energy landscape of the molecule based on two conventional backbone dihedral angles and . Figure 4 presents the free energy landscape as computed using traditional umbrella sampling, and superimposes a scatter plot of the samples gathered during the active learning loop of SBC. Note that the active learning loop is bootstrapped from equilibrium samples only, i.e. the model does not a priori know what the transition regions look like. Nevertheless, iterative application of harmonic bias potentials centered on the learned log probabilities (see Figure 1) yields a complete sampling of the phace space. In the Supplemental Material (Section S5), we further demonstrate that the learned log probabilities can be used to perform 1D umbrella sampling in order to estimate the relative stability between c7eq and c7ax, which is much more efficient than having to perform 2D umbrella sampling across the entire () domain.
By embedding SBC into the learned feature space of a GNN, we exploit the intrinsic overlap that exists between a quantitative understanding of atomic interactions and free energy minima that arise as a result of those interactions. To demonstrate the efficiency of this deliberate design choice, Table 1 enumerates and compares the most prominent CV learning methods with SBC for alanine dipeptide. We evaluate a number of key characteristics, including the required amount of input data to train a given CV, its transferability towards symmetrically equivalent geometries, whether or not it requires manual feature selection, and whether or not an initial transition path is required. In essence, Table 1 shows that SBC outperforms every single method in all of these aspects. It requires the fewest structures and smallest total simulation time; it is the only method that is rigorously invariant with respect to all symmetry operations; and it does not require any kind of manual input feature selection or initial transition path. While not highlighted in Table 1, it should also be noted that SBC is competitive in terms of its computational cost during inference, i.e. during production (enhanced sampling) simulations. Evaluation of the log probability difference as a learned CV during biased simulations is essentially trivial since the node features have already been computed and used for the (unbiased) energy and forces.
| training cost | input descriptors | symmetries | initialized without prior transition path(s) | ||||
| snapshots | time [ns] | trans | rot | perm | |||
| time-lagged AE [12] | 750 ns | learned111based on alignment with respect to a reference atomic geometry | |||||
| FEBILAE [27] | 1500 ns | learned1 | |||||
| VAMPnets [13] | 250,000 | 250 ns | learned1 | ✓ | |||
| LINES [28] | 100 ns | manual | |||||
| LED [29] | 100 ns | manual | |||||
| Ref. [30] | 300,000 | learned1 | ✓ | ||||
| Deep-LDA [9] | 10,000 | manual | |||||
| Deep-TDA [22] | 40 ns | learned222based on a list of interatomic distances | ✓ | ||||
| Ref. [31] | 15,000 | 15 ns | manual | ✓ | |||
| Ref. [32] | 8,000 | 512 ns | manual | ||||
| SBC | 3,694 | 11 ns | learned | ✓ | |||
We conclude with a discussion on the current limitations of SBC as well as directions for future research. Clearly, the success of our method is critically dependent on the availability of ML potentials which can predict the atomic interactions in the system with high fidelity. With the introduction of equivariant feature representations, specialized interaction layers, and efficient online learning workflows, such potentials can now be routinely developed without requiring massive amounts of QM reference data, even for challenging systems [16, 33, 34, 35]. In addition, the availability of foundation models – which are applicable to essentially the entire periodic table – can alleviate some of the burden that comes with training and validating the base GNN [36, 37, 38]. In this work, we considered both self-developed and foundation models as base GNN in which the SBC model is inserted – the alanine dipeptide and MIL-53(Al) case studies were performed with custom ML potentials trained using psiflow (see the Supplemental Material, Section S4), while a MACE foundation model was used for the proton transfer reaction. As such, we can conclude that SBC is fundamentally compatible with both. However, more challenging scenarios might require multi-task training of the entire GNN (including its interaction layers), in which each sample in the training set is labeled with QM reference data and additionally a phase or basin label – this will incentivize the model to design the per-atom feature embedding to be informative for both potential energy prediction as well as basin classification. Finally, we mention that the proposed method can be extended towards the more general case of free energy minima simply by increasing the number of classes in the classification layer. This will be necessary when considering more extreme scenarios such as large chemical reaction networks or protein conformational landscapes.
Acknowledgements.
The authors acknowledge funding from the Research Board of Ghent University (V.V.S) and iBOF-21-085 PERSIST (V.V.S. and T.B.). S.V. and P.D. wish to thank the Research Foundation – Flanders (FWO) for doctoral fellowships (grant nos. 11H6821N and 11O2123N respectively). M.B. acknowledges financial support from the Fund for Scientific Research Flanders and the Excellence of Science (EOS) Project BioFact (EOS ID 30902231). We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call. Part of the simulations were performed on the Luxembourg national supercomputer MeluXina. The authors gratefully acknowledge the LuxProvide teams for their expert support.References
- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications (Elsevier Science, 2023).
- Tuckerman [2010] M. Tuckerman, Statistical Mechanics: Theory and Molecular Simulation, Oxford Graduate Texts (OUP Oxford, 2010).
- Grifoni et al. [2019] E. Grifoni, G.-M. Piccini, and M. Parrinello, Microscopic description of acid–base equilibrium, Proc. Natl. Acad. Sci. U.S.A. 116, 4054 (2019).
- Ahalawat and Mondal [2018] N. Ahalawat and J. Mondal, Assessment and optimization of collective variables for protein conformational landscape: GB1 -hairpin as a case study, J. Chem. Phys. 149, 094101 (2018).
- Niu et al. [2018] H. Niu, P. M. Piaggi, M. Invernizzi, and M. Parrinello, Molecular dynamics simulations of liquid silica crystallization, Proc. Natl. Acad. Sci. U.S.A. 115, 5348 (2018).
- Chen [2021] M. Chen, Collective variable-based enhanced sampling and machine learning, Eur. Phys. J. B 94, 211 (2021).
- Bhakat [2022] S. Bhakat, Collective variable discovery in the age of machine learning: reality, hype and everything in between, RSC Adv. 12, 25010 (2022).
- Elishav et al. [2023] O. Elishav, R. Podgaetsky, O. Meikler, and B. Hirshberg, Collective variables for conformational polymorphism in molecular crystals, J. Phys. Chem. Lett. 14, 971 (2023).
- Bonati et al. [2020] L. Bonati, V. Rizzi, and M. Parrinello, Data-driven collective variables for enhanced sampling, J. Phys. Chem. Lett. 11, 2998 (2020).
- Mendels et al. [2018] D. Mendels, G. Piccini, and M. Parrinello, Collective variables from local fluctuations, J. Phys. Chem. Lett. 9, 2776 (2018).
- Pérez-Hernández et al. [2013] G. Pérez-Hernández, F. Paul, T. Giorgino, G. De Fabritiis, and F. Noé, Identification of slow molecular order parameters for markov model construction, J. Chem. Phys. 139, 015102 (2013).
- Wehmeyer and Noé [2018] C. Wehmeyer and F. Noé, Time-lagged autoencoders: Deep learning of slow collective variables for molecular kinetics, J. Chem. Phys. 148, 241703 (2018).
- Mardt et al. [2018] A. Mardt, L. Pasquali, H. Wu, and F. Noé, Vampnets for deep learning of molecular kinetics, Nat. Commun. 9, 5 (2018).
- Thomas et al. [2018] N. Thomas, T. Smidt, S. Kearnes, L. Yang, L. Li, K. Kohlhoff, and P. Riley, Tensor field networks: Rotation- and translation-equivariant neural networks for 3d point clouds (2018), arXiv:1802.08219 [cs.LG] .
- Batzner et al. [2022] S. Batzner, A. Musaelian, L. Sun, M. Geiger, J. P. Mailoa, M. Kornbluth, N. Molinari, T. E. Smidt, and B. Kozinsky, E(3)-equivariant graph neural networks for data-efficient and accurate interatomic potentials, Nat. Commun. 13, 2453 (2022).
- Kovács et al. [2023] D. P. Kovács, I. Batatia, E. S. Arany, and G. Csányi, Evaluation of the MACE force field architecture: From medicinal chemistry to materials science, J. Chem. Phys. 159, 044118 (2023).
- Bishop [2006] C. M. Bishop, Pattern Recognition and Machine Learning (Information Science and Statistics) (Springer-Verlag, Berlin, Heidelberg, 2006).
- Shirts and Chodera [2008] M. R. Shirts and J. D. Chodera, Statistically optimal analysis of samples from multiple equilibrium states, J. Chem. Phys. 129, 124105 (2008).
- Shirts [2017] M. R. Shirts, Reweighting from the mixture distribution as a better way to describe the multistate bennett acceptance ratio (2017), arXiv:1704.00891 [cond-mat.stat-mech] .
- Shirts and Ferguson [2020] M. R. Shirts and A. L. Ferguson, Statistically optimal continuous free energy surfaces from biased simulations and multistate reweighting, J. Chem. Theory Comput. 16, 4107 (2020).
- Shirts et al. [2023] M. Shirts, K. Beauchamp, L. Naden, J. Chodera, jaimergp, S. Martiniani, C. Stern, M. Henry, J. Fass, R. Gowers, R. T. McGibbon, B. Dice, C. Jones, D. L. Dotson, F. York, I. Pulido, M. Thompson, and T. Burgin, choderalab/pymbar: 4.0.2 optional jax build (2023).
- Trizio and Parrinello [2021] E. Trizio and M. Parrinello, From enhanced sampling to reaction profiles, The Journal of Physical Chemistry Letters 12, 8621 (2021).
- Serre et al. [2002] C. Serre, F. Millange, C. Thouvenot, M. Noguès, G. Marsolier, D. Louër, and G. Férey, Very large breathing effect in the first nanoporous chromium(iii)-based solids: MIL-53 or CrIII(OH)·{O2C-C6H4-CO2}·{HO2C-C6H4-CO2H}x·H2Oy, Journal of the American Chemical Society 124, 13519 (2002).
- Demuynck et al. [2018] R. Demuynck, J. Wieme, S. M. J. Rogge, K. D. Dedecker, L. Vanduyfhuys, M. Waroquier, and V. Van Speybroeck, Protocol for identifying accurate collective variables in enhanced molecular dynamics simulations for the description of structural transformations in flexible metal–organic frameworks, Journal of Chemical Theory and Computation 14, 5511 (2018).
- Tan et al. [2024] A. R. Tan, J. C. B. Dietschreit, and R. Gomez-Bombarelli, Enhanced sampling of robust molecular datasets with uncertainty-based collective variables (2024), arXiv:2402.03753 [cs.LG] .
- Bailleul et al. [2020] S. Bailleul, K. Dedecker, P. Cnudde, L. Vanduyfhuys, M. Waroquier, and V. Van Speybroeck, Ab initio enhanced sampling kinetic study on mto ethene methylation reaction, Journal of Catalysis 388, 38 (2020).
- Belkacemi et al. [2022] Z. Belkacemi, P. Gkeka, T. Lelièvre, and G. Stoltz, Chasing collective variables using autoencoders and biased trajectories, J. Chem. Theory Comput. 18, 59 (2022).
- Odstrcil et al. [2022] R. E. Odstrcil, P. Dutta, and J. Liu, Lines: Log-probability estimation via invertible neural networks for enhanced sampling, J. Chem. Theory Comput. 18, 6297 (2022).
- Vlachas et al. [2022] P. R. Vlachas, J. Zavadlav, M. Praprotnik, and P. Koumoutsakos, Accelerated simulations of molecular systems through learning of effective dynamics, J. Chem. Theory Comput. 18, 538 (2022).
- Sun et al. [2022] L. Sun, J. Vandermause, S. Batzner, Y. Xie, D. Clark, W. Chen, and B. Kozinsky, Multitask machine learning of collective variables for enhanced sampling of rare events, J. Chem. Theory Comput. 18, 2341 (2022).
- Bonati et al. [2021] L. Bonati, G. Piccini, and M. Parrinello, Deep learning the slow modes for rare events sampling, Proc. Natl. Acad. Sci. U.S.A. 118, e2113533118 (2021).
- Müllender et al. [2023] L. Müllender, A. Rizzi, M. Parrinello, P. Carloni, and D. Mandelli, Effective data-driven collective variables for free energy calculations from metadynamics of paths (2023), arXiv:2311.05571 [physics.comp-ph] .
- Vandenhaute et al. [2023] S. Vandenhaute, M. Cools-Ceuppens, S. DeKeyser, T. Verstraelen, and V. Van Speybroeck, Machine learning potentials for metal-organic frameworks using an incremental learning approach, npj Comput. Mater. 9, 19 (2023).
- Liao and Smidt [2023] Y.-L. Liao and T. Smidt, Equiformer: Equivariant graph attention transformer for 3d atomistic graphs (2023), arXiv:2206.11990 [cs.LG] .
- Pelaez et al. [2024] R. P. Pelaez, G. Simeon, R. Galvelis, A. Mirarchi, P. Eastman, S. Doerr, P. Thölke, T. E. Markland, and G. D. Fabritiis, Torchmd-net 2.0: Fast neural network potentials for molecular simulations (2024), arXiv:2402.17660 [cs.LG] .
- Batatia et al. [2023] I. Batatia, P. Benner, Y. Chiang, A. M. Elena, D. P. Kovács, J. Riebesell, X. R. Advincula, M. Asta, W. J. Baldwin, N. Bernstein, A. Bhowmik, S. M. Blau, V. Cărare, J. P. Darby, S. De, F. D. Pia, V. L. Deringer, R. Elijošius, Z. El-Machachi, E. Fako, A. C. Ferrari, A. Genreith-Schriever, J. George, R. E. A. Goodall, C. P. Grey, S. Han, W. Handley, H. H. Heenen, K. Hermansson, C. Holm, J. Jaafar, S. Hofmann, K. S. Jakob, H. Jung, V. Kapil, A. D. Kaplan, N. Karimitari, N. Kroupa, J. Kullgren, M. C. Kuner, D. Kuryla, G. Liepuoniute, J. T. Margraf, I.-B. Magdău, A. Michaelides, J. H. Moore, A. A. Naik, S. P. Niblett, S. W. Norwood, N. O’Neill, C. Ortner, K. A. Persson, K. Reuter, A. S. Rosen, L. L. Schaaf, C. Schran, E. Sivonxay, T. K. Stenczel, V. Svahn, C. Sutton, C. van der Oord, E. Varga-Umbrich, T. Vegge, M. Vondrák, Y. Wang, W. C. Witt, F. Zills, and G. Csányi, A foundation model for atomistic materials chemistry (2023), arXiv:2401.00096 [physics.chem-ph] .
- Deng et al. [2023] B. Deng, P. Zhong, K. Jun, J. Riebesell, K. Han, C. J. Bartel, and G. Ceder, Chgnet as a pretrained universal neural network potential for charge-informed atomistic modelling, Nature Machine Intelligence 5, 1031 (2023).
- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).

### v3

@@
-Free Energy Calculations using Smooth Basin Classification
+Rare Event Sampling using Smooth Basin Classification
 Abstract
@@
 While explicit integration of Newton’s equation of motion requires timesteps in the order of a femtosecond, the time required for many transformations to complete is often beyond the micro- or millisecond range because of large energy barriers between initial and final state.
-As such, simulation of a single transformation can easily require billions of time steps [1, 2].
+Such transformations are therefore almost always rare events, and their simulation can easily require billions of time steps [1, 2].
 Over the past decades, a variety of advanced sampling techniques have been proposed in order to improve the sampling efficiency of dynamic simulations.
@@
 The bias energy is defined along a certain predefined reaction coordinate or collective variable (CV).
-The CV is a function of the atomic coordinates and should naturally discriminate between initial and final state without violating the physical symmetries in the system.
+The CV is a function of the atomic coordinates and should naturally discriminate between initial and final state without violating the physical symmetries in the system [3].
 Specifically, the CV should remain invariant with respect to global translations and rotations, as well as permutations of atoms which are chemically identical.
@@
 Considering the numerous achievements of machine learning in atomistic modeling, it is sensible to anticipate that suitable reaction coordinates can also be learned from (unbiased) simulation data, and ideally without requiring human intuition about the transition mechanism(s).
-A wide variety of such methods have been proposed over the past decade, and we briefly mention a few notable examples while referring the reader to specialized reviews for a more comprehensive discussion [6, 7].
-Supervised methods learn a continuous discriminant function which separates free energy minima based on user-defined structural descriptors, for example using (deep) linear discriminant analysis [8, 9, 10].
+A wide variety of such methods have been proposed over the past decade, and we briefly mention a few notable examples while referring the reader to specialized reviews for a more comprehensive discussion [7, 8, 9].
+Supervised methods learn a continuous discriminant function which separates free energy minima based on user-defined structural descriptors, for example using (deep) linear discriminant analysis [10, 11, 12, 13] or (variational) autoencoders [14, 15, 16].
 Unsupervised methods do not directly learn a discriminant function but instead identify so-called slow modes, either within a given set of descriptors or entirely end-to-end; see e.g.
@@
 In this work, we propose smooth basin classification (SBC); a universal method to construct CVs and compute free energy differences for chemical and physical transformations.
-SBC builds upon the successful development of graph neural networks (GNNs) as effective interatomic potentials by using their learned feature space as ansatz for constructing physically meaningful CVs [14, 15, 16].
-We apply SBC for complex rare events in both materials science and chemistry, and demonstrate that it matches or outperforms CVs based on human intuition.
-Relative free energy predictions based on learned CVs are within 1 kJ/mol of reference values as computed based on established CVs for each of the case studies.
-In addition, we observe that SBC requires orders of magnitude less training data as compared to existing approaches [6, 7, 8, 9, 10, 11, 12, 13].
-This is a direct consequence of the fact that the constructed CVs are embedded in the learned feature space of the GNN, thereby inheriting all of their established benefits regarding data efficiency and universality.
+SBC builds upon the successful development of graph neural networks (GNNs) as effective interatomic potentials by using their learned feature space as ansatz for constructing physically meaningful CVs [20, 21, 22, 23].
+Our key finding is that GNN-based interaction potentials produce a feature representation for atomic environments which can be used to drive transitions between free energy minima and characterize their relative stability.
 Consider a system with two basins of attraction, named A and B.
@@
 Instead of defining this mapping directly in terms of the atomic coordinates – which are not invariant with respect to the symmetry operations in the system – we embed it into the learned feature space of a GNN which has been pretrained to reproduce the potential energy surface of the system (Figure 1, b).
-In general, the interaction layers in the GNN parse the chemical environment of atom into a feature vector in this space (which typically contains anywhere between 8 and 64 dimensions).
-Importantly, this feature vector is rigorously invariant with respect to all symmetry operations in the system, and is specifically trained to capture the physical interactions in the system.
-Both the atomic energy and the total energy are computed from the atomic feature vectors using a simple readout function:
-Because the atomic features are maximally informative for determining the interaction energy of a given atomic environment, we may assume that they are also highly informative in predicting the basin of attraction, i.e.
-the final destination of an energy minimization.
-In full analogy with equation 1, we can therefore construct a prediction for the basin of attraction using the same GNN but with a different readout function (Figure 1, c):
+In general, the interaction layers of the GNN parse the chemical environment of atom into a feature vector in this space.
+Importantly, this feature vector is rigorously invariant with respect to all symmetry operations in the system, and is specifically trained to describe the atomic interactions.
+Both the per-atom energy and the total energy are computed from the feature vectors using a simple readout function:
+in which denotes the total number of atoms in configuration .
+Because the features are maximally informative for determining the interaction energy of a given atomic environment, we may assume that they are also highly informative in predicting the basin of attraction.
+In full analogy with equation 1, we can augment the GNN with an additional classification readout function (Figure 1, c):
 in which we introduce the per-atom log probabilities as an invariant tuple which measures how strongly the model associates a given atomic environment to basin A or B – as commonly done in a classification setting [24]:
 For example, if the model believes that a given atomic environment is more likely to appear in phase as opposed to phase , it will predict log probabilities such that .
-After summing the per-atom log probabilities to obtain the total log probabilities , the predicted basin of attraction of the total configuration is then determined by the phase which received the largest log probability in :
-A more detailed overview of the model architecture can be found in the Supplemental Material (Section S1).
-While the prediction of the basin of attraction is essentially a discrete classification task, the actual log probabilities are smooth functions of the atomic feature vectors and hence the atomic coordinates – which means that they can be used to introduce biasing forces and/or virial stress.
+After summing the per-atom log probabilities to obtain the total log probabilities , the predicted basin of attraction of the total configuration is then determined by the largest component in .
+A detailed overview of the phase classification readout can be found in the Appendix.
+While the prediction of the basin of attraction is essentially a discrete classification task, the actual log probabilities are smooth functions of the atomic feature vectors and hence the atomic coordinates – which means that they can be used to introduce biasing forces (and biasing stress).
 Ideally, we would like to design this bias such that it can drive transitions from and to any of the basins.
@@
 It is therefore strictly positive in basin B and strictly negative in basin A provided that the phase readout is properly trained (Figure 1, d).
-Clearly, the quantity correctly discriminates between two free energy minima while also being continuously differentiable with respect to the atomic coordinates (using standard backpropagation).
+Clearly, the quantity correctly discriminates between two basins of attraction while also being continuously differentiable with respect to the atomic coordinates (using standard backpropagation).
 It is therefore, by definition, a valid collective variable for the transition between basins A and B.
-Note that the invariance of with respect to global translations and rotations as well as permutation of chemically identical atoms is guaranteed by construction of the GNN.
-The built-in symmetry of SBC towards translations, rotations, and permutations incorporates important physical behavior into the model, and should be contrasted with both common practice CVs and most other ML-based approaches [9, 13, 12, 22].
-which often use functions and/or input features that depend explicitly on the atomic indices (thereby violating permutation invariance) or specific components of the unit cell vectors (thereby violating rotational as well as periodic invariance).
-The phase readout layers are analogous to the existing energy readouts found in conventional GNNs [14, 15, 16] (Figure 1, c), and essentially consist of a multilayer perceptron (MLP) containing at most two nonlinearities and about 1000 weights in total (compared to 100,000 weights in the GNN).
-The classifier is trained while keeping the GNN and energy readouts fixed, similar to conventional transfer learning approaches.
-Its training data consists of atomic geometries which are labeled with their corresponding basin of attraction (e.g.
-A or B).
-To ensure that the classifier remains accurate for out-of-equilibrium geometries, we employ a combination of enhanced sampling (in ) and active learning to explore transition regions between phases in an iterative fashion (Figure 1, e).
-More information regarding model training and the active learning approach can be found in the Supplemental Material (Sections S1 and S2).
+In what follows, we discuss the application of SBC for prototypical physical and chemical transformations.
+We use MACE as base GNN [25, 22], and employ active learning as in Figure 1 (e) to train the classifier readout – see the appendix for additional computational details.
 Figure 2 (a) shows a nonreactive phase transition in a solid-state system named MIL-53(Al), a flexible metal-organic framework based on a winerack topology [26].
-The system exhibits two basins of attraction which are relevant for tuning the material towards gas storage and/or sensing applications; a closed pore () and open pore () phase.
+The system exhibits two basins of attraction which are relevant for tuning the material towards gas storage and/or sensing applications; a closed pore () and open pore () phase [27].
 In these types of transitions, i.e.
@@
 Figure 2 (a) demonstrates the performance of the unit cell volume as formal discriminator between the two basins, and compares it with a difference of learned log probabilities as proposed in equation 4.
-Both criteria perform excellent in terms of classification accuracy, and in addition exhibit a very strong correlation.
-For both collective variables, the misclassified geometries are all localized near the transition region (such they will not cause substantial differences in the relative stability).
+The misclassified geometries are all localized near the transition region (such they will not cause substantial differences in the relative stability).
 Note that the network learns to distinguish low- and high-volume configurations based on atomic environments only, i.e.
 without explicitly relying on unit cell parameters as input.
-Using the log probability difference as CV (see equation 5), we can employ conventional enhanced sampling techniques and evaluate the free energy profile .
-Here, we choose to employ umbrella sampling with harmonic restraints on the CV:
-in which we emphasize that the bias energy contains an implicit dependence on both the atomic coordinates and the unit cell parameters through the learned atomic environment features [25].
-We use the multistate Bennett acceptance ratio (MBAR) equations to combine samples from multiple umbrellas in a statistically optimal manner and obtain free energy profiles as shown in Figure 2 (b, top) [18, 19, 20, 21].
-The reference free energy profile using the unit cell volume as bias coordinate is added for completeness in Figure 2 (b, bottom).
-The total relative free energy difference between the and basin, as obtained by integrating each of the profiles in their respective minima, agrees to within 1 kJ/mol between both CVs.
-Note that while relative stabilities are intrinsic to the system, the actual free energy profiles depend on the particular choice of CV, so the learned and reference free energy profiles need not have the same shape (although there exist ways to transform one profile into another [26]).
-For global phase transformations in solids such as the one in Figure 2, a size-extensive definition of as a sum over per-atom log probabilities is intuitively sensible because it gives the model the ability to assign nonzero log probability contributions to all atoms simultaneously.
+| training cost | input descriptors | symmetries | initialized without prior transition path(s) | ||||
+| snapshots | time [ns] | trans | rot | perm | |||
+| time-lagged AE [18] | 750 ns | learned | |||||
+| FEBILAE [29] | 1500 ns | learned | |||||
+| VAMPnets [19] | 250,000 | 250 ns | learned | ✓ | |||
+| LINES [30] | 100 ns | manual | |||||
+| LED [31] | 100 ns | manual | |||||
+| Ref. [32] | 300,000 | learned | ✓ | ||||
+| Deep-LDA [11] | 10,000 | manual | |||||
+| Deep-TDA [33] | 40 ns | learned | ✓ | ||||
+| Ref. [34] | 15,000 | 15 ns | manual | ✓ | |||
+| Ref. [35] | 8,000 | 512 ns | manual | ||||
+| SBC | 3,694 | 11 ns | learned | ✓ | |||
+For global phase transformations in solids such as the one in Figure 2 (a), a size-extensive definition of as a sum over per-atom log probabilities is intuitively sensible because it gives the model the ability to assign nonzero log probability contributions to all atoms simultaneously.
 However, in chemical transformations, the rare event is most often a local process whose free energy barrier does not necessarily depend on the size of the surrounding environment (e.g.
 the catalyst).
-To demonstrate the applicability of our approach to those cases, we investigate a proton transfer reaction from a Brønsted acid site in a zeolite to an alkene guest (Figure 3, a).
-The reference collective variable in this case is a coordination number difference which is defined explicitly in terms of the indices of the oxygen atoms of the active site as well as the indices of the carbon atoms of the molecule and all of the hydrogens – see the Supplemental Material (Section S5) for more details.
-While such definitions are extremely common, they violate permutation symmetry and make each collective variable highly system-specific.
-Figure 3 (a) shows the classification performance of the coordination number versus a learned CV based on equation 5.
-Again, classification accuracies are excellent for both the proposed CV based on chemical intuition as well as the learned log probability difference.
-Misclassified geometries are localized at or near the transition region, which means that they will not contribute to the relative free energy of both phases.
-Indeed, as indicated in Figure 3 (b), the free energy difference between both basins as computed from either of the two profiles agrees to within 1 kJ/mol.
-A final case study concerns the conformational free energy landscape of alanine dipeptide, given that it is the most commonly studied system in the context of free energy calculations and the automated development of collective variables [12, 27, 13, 28, 29, 30, 9, 31, 32].
-At the PBE-D3(BJ) level of theory and room temperature conditions, the molecule is observed to exhibit two main stable conformations; the c7eq and c7ax conformations.
-There exist multiple transition paths between both conformations, and each comprises multiple potential energy minima.
-The free energy of the system is most often visualized using a two-dimensional free energy landscape of the molecule based on two conventional backbone dihedral angles and .
-Figure 4 presents the free energy landscape as computed using traditional umbrella sampling, and superimposes a scatter plot of the samples gathered during the active learning loop of SBC.
-Note that the active learning loop is bootstrapped from equilibrium samples only, i.e.
-the model does not a priori know what the transition regions look like.
-Nevertheless, iterative application of harmonic bias potentials centered on the learned log probabilities (see Figure 1) yields a complete sampling of the phace space.
-In the Supplemental Material (Section S5), we further demonstrate that the learned log probabilities can be used to perform 1D umbrella sampling in order to estimate the relative stability between c7eq and c7ax, which is much more efficient than having to perform 2D umbrella sampling across the entire () domain.
-By embedding SBC into the learned feature space of a GNN, we exploit the intrinsic overlap that exists between a quantitative understanding of atomic interactions and free energy minima that arise as a result of those interactions.
-To demonstrate the efficiency of this deliberate design choice, Table 1 enumerates and compares the most prominent CV learning methods with SBC for alanine dipeptide.
+To demonstrate the applicability of our approach to those cases, we investigate a proton transfer reaction from a Brønsted acid site in a zeolite to an alkene guest (Figure 2, b).
+The reference collective variable in this case is a coordination number difference which is defined explicitly in terms of the indices of the oxygen atoms of the active site as well as the indices of the carbon atoms of the molecule and all of the hydrogens – see the Supplemental Material (Section S1.2) for more details.
+Classification accuracies are excellent for both the proposed CV based on chemical intuition as well as the learned log probability difference.
+Misclassified geometries are localized at or near the transition region, which means that they will not induce discrepancies in the relative stability.
+Indeed, for both systems, free energy calculations based on either the reference or the learned CV yield the same relative stability (to within 1 kJ/mol).
+Note that the quantity satisfies the same symmetries as the potential energy of the system (Figure 1, b/c).
+As such, the learned collective variable is invariant with respect to global translations and rotations as well as permutations of chemically identical atoms.
+Achieving permutational symmetry in particular is quite remarkable, and should be contrasted with both common practice CVs and most other ML-based approaches [11, 19, 18, 33] which often use functions and/or input features that depend explicitly on the atomic indices (thereby violating permutation invariance) or specific components of the unit cell vectors (thereby violating rotational as well as periodic invariance).
+We hypothesize that the Euclidean and permutation symmetry of is the primary reason for the exceptional data efficiency of the method, alongside the effective active learning scheme (Figure 1, e) and the implicit use of equivariant features within the GNN.
+We demonstrate this explicitly based on a case study on alanine dipeptide, which represents a topical nonreactive conformational change that has served as enhanced sampling benchmark system in many studies [18, 29, 19, 30, 31, 32, 11, 34, 35, 36].
+Table 1 enumerates and compares the most prominent CV learning methods with SBC for alanine dipeptide.
 We evaluate a number of key characteristics, including the required amount of input data to train a given CV, its transferability towards symmetrically equivalent geometries, whether or not it requires manual feature selection, and whether or not an initial transition path is required.
@@
 It requires the fewest structures and smallest total simulation time; it is the only method that is rigorously invariant with respect to all symmetry operations; and it does not require any kind of manual input feature selection or initial transition path.
-While not highlighted in Table 1, it should also be noted that SBC is competitive in terms of its computational cost during inference, i.e.
-during production (enhanced sampling) simulations.
-Evaluation of the log probability difference as a learned CV during biased simulations is essentially trivial since the node features have already been computed and used for the (unbiased) energy and forces.
-| training cost | input descriptors | symmetries | initialized without prior transition path(s) | ||||
-| snapshots | time [ns] | trans | rot | perm | |||
-| time-lagged AE [12] | 750 ns | learned111based on alignment with respect to a reference atomic geometry | |||||
-| FEBILAE [27] | 1500 ns | learned1 | |||||
-| VAMPnets [13] | 250,000 | 250 ns | learned1 | ✓ | |||
-| LINES [28] | 100 ns | manual | |||||
-| LED [29] | 100 ns | manual | |||||
-| Ref. [30] | 300,000 | learned1 | ✓ | ||||
-| Deep-LDA [9] | 10,000 | manual | |||||
-| Deep-TDA [22] | 40 ns | learned222based on a list of interatomic distances | ✓ | ||||
-| Ref. [31] | 15,000 | 15 ns | manual | ✓ | |||
-| Ref. [32] | 8,000 | 512 ns | manual | ||||
-| SBC | 3,694 | 11 ns | learned | ✓ | |||
+Figure 3 demonstrates how SBC is capable of discovering several transition paths during its active learning loop.
+In Section S1.3 of the Supplemental Material, we further show that the relative free energy difference as computed using and agrees within 1 kJ/mol.
 We conclude with a discussion on the current limitations of SBC as well as directions for future research.
 Clearly, the success of our method is critically dependent on the availability of ML potentials which can predict the atomic interactions in the system with high fidelity.
-With the introduction of equivariant feature representations, specialized interaction layers, and efficient online learning workflows, such potentials can now be routinely developed without requiring massive amounts of QM reference data, even for challenging systems [16, 33, 34, 35].
-In addition, the availability of foundation models – which are applicable to essentially the entire periodic table – can alleviate some of the burden that comes with training and validating the base GNN [36, 37, 38].
-In this work, we considered both self-developed and foundation models as base GNN in which the SBC model is inserted – the alanine dipeptide and MIL-53(Al) case studies were performed with custom ML potentials trained using psiflow (see the Supplemental Material, Section S4), while a MACE foundation model was used for the proton transfer reaction.
+In this work, we considered both self-developed and foundation models as base GNN in which the SBC model is inserted – the alanine dipeptide and MIL-53(Al) case studies were performed with MACE models trained using psiflow [22, 37], while a MACE foundation model was used for the proton transfer reaction[38].
 As such, we can conclude that SBC is fundamentally compatible with both.
 However, more challenging scenarios might require multi-task training of the entire GNN (including its interaction layers), in which each sample in the training set is labeled with QM reference data and additionally a phase or basin label – this will incentivize the model to design the per-atom feature embedding to be informative for both potential energy prediction as well as basin classification.
+In addition to relative free energies, the learned CVs can be employed to predict the kinetics of rare events based on e.g.
+reactive flux and/or path sampling methods [39].
 Finally, we mention that the proposed method can be extended towards the more general case of free energy minima simply by increasing the number of classes in the classification layer.
-This will be necessary when considering more extreme scenarios such as large chemical reaction networks or protein conformational landscapes.
-Acknowledgements.
+DATA AVAILABILITY STATEMENT
+An implementation of SBC in MACE will be made available in a public GitHub repository upon final publication; for peer-review purposes, we also provide the code through the following link: https://tinyurl.com/mufvj4ur.
+All additional input files, all generated datasets, and all of the models which were trained in this manuscript will be made available in a public Zenodo archive upon final publication; for peer-review purposes, these are available upon request as well.
+ACKNOWLEDGEMENTS
 The authors acknowledge funding from the Research Board of Ghent University (V.V.S) and iBOF-21-085 PERSIST (V.V.S.
@@
 Part of the simulations were performed on the Luxembourg national supercomputer MeluXina.
-The authors gratefully acknowledge the LuxProvide teams for their expert support.References
-- Frenkel and Smit [2023] D. Frenkel and B. Smit, Understanding Molecular Simulation: From Algorithms to Applications (Elsevier Science, 2023).
-- Tuckerman [2010] M. Tuckerman, Statistical Mechanics: Theory and Molecular Simulation, Oxford Graduate Texts (OUP Oxford, 2010).
-- Grifoni et al. [2019] E. Grifoni, G.-M. Piccini, and M. Parrinello, Microscopic description of acid–base equilibrium, Proc. Natl. Acad. Sci. U.S.A. 116, 4054 (2019).
-- Ahalawat and Mondal [2018] N.
-Ahalawat and J.
-Mondal, Assessment and optimization of collective variables for protein conformational landscape: GB1 -hairpin as a case study, J.
-Chem.
-Phys.
-149, 094101 (2018).
-- Niu et al. [2018] H. Niu, P. M. Piaggi, M. Invernizzi, and M. Parrinello, Molecular dynamics simulations of liquid silica crystallization, Proc. Natl. Acad. Sci. U.S.A. 115, 5348 (2018).
-- Chen [2021] M. Chen, Collective variable-based enhanced sampling and machine learning, Eur. Phys. J. B 94, 211 (2021).
-- Bhakat [2022] S. Bhakat, Collective variable discovery in the age of machine learning: reality, hype and everything in between, RSC Adv. 12, 25010 (2022).
-- Elishav et al. [2023] O. Elishav, R. Podgaetsky, O. Meikler, and B. Hirshberg, Collective variables for conformational polymorphism in molecular crystals, J. Phys. Chem. Lett. 14, 971 (2023).
-- Bonati et al. [2020] L. Bonati, V. Rizzi, and M. Parrinello, Data-driven collective variables for enhanced sampling, J. Phys. Chem. Lett. 11, 2998 (2020).
-- Mendels et al. [2018] D. Mendels, G. Piccini, and M. Parrinello, Collective variables from local fluctuations, J. Phys. Chem. Lett. 9, 2776 (2018).
-- Pérez-Hernández et al.
-[2013] G.
-Pérez-Hernández, F.
-Paul, T.
-Giorgino, G.
-De Fabritiis, and F.
-Noé, Identification of slow molecular order parameters for markov model construction, J.
-Chem.
-Phys.
-139, 015102 (2013).
-- Wehmeyer and Noé [2018] C. Wehmeyer and F. Noé, Time-lagged autoencoders: Deep learning of slow collective variables for molecular kinetics, J. Chem. Phys. 148, 241703 (2018).
-- Mardt et al. [2018] A. Mardt, L. Pasquali, H. Wu, and F. Noé, Vampnets for deep learning of molecular kinetics, Nat. Commun. 9, 5 (2018).
-- Thomas et al.
-[2018] N.
-Thomas, T.
-Smidt, S.
-Kearnes, L.
-Yang, L.
-Li, K.
-Kohlhoff, and P.
-Riley, Tensor field networks: Rotation- and translation-equivariant neural networks for 3d point clouds (2018), arXiv:1802.08219 [cs.LG] .
-- Batzner et al.
-[2022] S.
-Batzner, A.
-Musaelian, L.
-Sun, M.
-Geiger, J.
-P.
-Mailoa, M.
-Kornbluth, N.
-Molinari, T.
-E.
-Smidt, and B.
-Kozinsky, E(3)-equivariant graph neural networks for data-efficient and accurate interatomic potentials, Nat.
-Commun.
-13, 2453 (2022).
-- Kovács et al.
-[2023] D.
-P.
-Kovács, I.
-Batatia, E.
-S.
-Arany, and G.
-Csányi, Evaluation of the MACE force field architecture: From medicinal chemistry to materials science, J.
-Chem.
-Phys.
-159, 044118 (2023).
-- Bishop [2006] C. M. Bishop, Pattern Recognition and Machine Learning (Information Science and Statistics) (Springer-Verlag, Berlin, Heidelberg, 2006).
-- Shirts and Chodera [2008] M. R. Shirts and J. D. Chodera, Statistically optimal analysis of samples from multiple equilibrium states, J. Chem. Phys. 129, 124105 (2008).
-- Shirts [2017] M. R. Shirts, Reweighting from the mixture distribution as a better way to describe the multistate bennett acceptance ratio (2017), arXiv:1704.00891 [cond-mat.stat-mech] .
-- Shirts and Ferguson [2020] M.
-R.
-Shirts and A.
-L.
-Ferguson, Statistically optimal continuous free energy surfaces from biased simulations and multistate reweighting, J.
-Chem.
-Theory Comput.
-16, 4107 (2020).
-- Shirts et al.
-[2023] M.
-Shirts, K.
-Beauchamp, L.
-Naden, J.
-Chodera, jaimergp, S.
-Martiniani, C.
-Stern, M.
-Henry, J.
-Fass, R.
-Gowers, R.
-T.
-McGibbon, B.
-Dice, C.
-Jones, D.
-L.
-Dotson, F.
-York, I.
-Pulido, M.
-Thompson, and T.
-Burgin, choderalab/pymbar: 4.0.2 optional jax build (2023).
-- Trizio and Parrinello [2021] E. Trizio and M. Parrinello, From enhanced sampling to reaction profiles, The Journal of Physical Chemistry Letters 12, 8621 (2021).
-- Serre et al.
-[2002] C.
-Serre, F.
-Millange, C.
-Thouvenot, M.
-Noguès, G.
-Marsolier, D.
-Louër, and G.
-Férey, Very large breathing effect in the first nanoporous chromium(iii)-based solids: MIL-53 or CrIII(OH)·{O2C-C6H4-CO2}·{HO2C-C6H4-CO2H}x·H2Oy, Journal of the American Chemical Society 124, 13519 (2002).
-- Demuynck et al.
-[2018] R.
-Demuynck, J.
-Wieme, S.
-M.
-J.
-Rogge, K.
-D.
-Dedecker, L.
-Vanduyfhuys, M.
-Waroquier, and V.
-Van Speybroeck, Protocol for identifying accurate collective variables in enhanced molecular dynamics simulations for the description of structural transformations in flexible metal–organic frameworks, Journal of Chemical Theory and Computation 14, 5511 (2018).
-- Tan et al.
-[2024] A.
-R.
-Tan, J.
-C.
-B.
-Dietschreit, and R.
-Gomez-Bombarelli, Enhanced sampling of robust molecular datasets with uncertainty-based collective variables (2024), arXiv:2402.03753 [cs.LG] .
-- Bailleul et al.
-[2020] S.
-Bailleul, K.
-Dedecker, P.
-Cnudde, L.
-Vanduyfhuys, M.
-Waroquier, and V.
-Van Speybroeck, Ab initio enhanced sampling kinetic study on mto ethene methylation reaction, Journal of Catalysis 388, 38 (2020).
-- Belkacemi et al. [2022] Z. Belkacemi, P. Gkeka, T. Lelièvre, and G. Stoltz, Chasing collective variables using autoencoders and biased trajectories, J. Chem. Theory Comput. 18, 59 (2022).
-- Odstrcil et al. [2022] R. E. Odstrcil, P. Dutta, and J. Liu, Lines: Log-probability estimation via invertible neural networks for enhanced sampling, J. Chem. Theory Comput. 18, 6297 (2022).
-- Vlachas et al.
-[2022] P.
-R.
-Vlachas, J.
-Zavadlav, M.
-Praprotnik, and P.
-Koumoutsakos, Accelerated simulations of molecular systems through learning of effective dynamics, J.
-Chem.
-Theory Comput.
-18, 538 (2022).
-- Sun et al.
-[2022] L.
-Sun, J.
-Vandermause, S.
-Batzner, Y.
-Xie, D.
-Clark, W.
-Chen, and B.
-Kozinsky, Multitask machine learning of collective variables for enhanced sampling of rare events, J.
-Chem.
-Theory Comput.
-18, 2341 (2022).
-- Bonati et al. [2021] L. Bonati, G. Piccini, and M. Parrinello, Deep learning the slow modes for rare events sampling, Proc. Natl. Acad. Sci. U.S.A. 118, e2113533118 (2021).
-- Müllender et al.
-[2023] L.
-Müllender, A.
-Rizzi, M.
-Parrinello, P.
-Carloni, and D.
-Mandelli, Effective data-driven collective variables for free energy calculations from metadynamics of paths (2023), arXiv:2311.05571 [physics.comp-ph] .
-- Vandenhaute et al.
-[2023] S.
-Vandenhaute, M.
-Cools-Ceuppens, S.
-DeKeyser, T.
-Verstraelen, and V.
-Van Speybroeck, Machine learning potentials for metal-organic frameworks using an incremental learning approach, npj Comput.
-Mater.
-9, 19 (2023).
-- Liao and Smidt [2023] Y.-L. Liao and T. Smidt, Equiformer: Equivariant graph attention transformer for 3d atomistic graphs (2023), arXiv:2206.11990 [cs.LG] .
-- Pelaez et al.
-[2024] R.
-P.
-Pelaez, G.
-Simeon, R.
-Galvelis, A.
-Mirarchi, P.
-Eastman, S.
-Doerr, P.
-Thölke, T.
-E.
-Markland, and G.
-D.
-Fabritiis, Torchmd-net 2.0: Fast neural network potentials for molecular simulations (2024), arXiv:2402.17660 [cs.LG] .
-- Batatia et al.
-[2023] I.
-Batatia, P.
-Benner, Y.
-Chiang, A.
-M.
-Elena, D.
-P.
-Kovács, J.
-Riebesell, X.
-R.
-Advincula, M.
-Asta, W.
-J.
-Baldwin, N.
-Bernstein, A.
-Bhowmik, S.
-M.
-Blau, V.
-Cărare, J.
-P.
-Darby, S.
-De, F.
-D.
-Pia, V.
-L.
-Deringer, R.
-Elijošius, Z.
-El-Machachi, E.
-Fako, A.
-C.
-Ferrari, A.
-Genreith-Schriever, J.
-George, R.
-E.
-A.
-Goodall, C.
-P.
-Grey, S.
-Han, W.
-Handley, H.
-H.
-Heenen, K.
-Hermansson, C.
-Holm, J.
-Jaafar, S.
-Hofmann, K.
-S.
-Jakob, H.
-Jung, V.
-Kapil, A.
-D.
-Kaplan, N.
-Karimitari, N.
-Kroupa, J.
-Kullgren, M.
-C.
-Kuner, D.
-Kuryla, G.
-Liepuoniute, J.
-T.
-Margraf, I.-B.
-Magdău, A.
-Michaelides, J.
-H.
-Moore, A.
-A.
-Naik, S.
-P.
-Niblett, S.
-W.
-Norwood, N.
-O’Neill, C.
-Ortner, K.
-A.
-Persson, K.
-Reuter, A.
-S.
-Rosen, L.
-L.
-Schaaf, C.
-Schran, E.
-Sivonxay, T.
-K.
-Stenczel, V.
-Svahn, C.
-Sutton, C.
-van der Oord, E.
-Varga-Umbrich, T.
-Vegge, M.
-Vondrák, Y.
-Wang, W.
-C.
-Witt, F.
-Zills, and G.
-Csányi, A foundation model for atomistic materials chemistry (2023), arXiv:2401.00096 [physics.chem-ph] .
-- Deng et al.
-[2023] B.
-Deng, P.
-Zhong, K.
-Jun, J.
-Riebesell, K.
-Han, C.
-J.
-Bartel, and G.
-Ceder, Chgnet as a pretrained universal neural network potential for charge-informed atomistic modelling, Nature Machine Intelligence 5, 1031 (2023).
-- Chen and Ong [2022] C. Chen and S. P. Ong, A universal graph deep learning interatomic potential for the periodic table, Nature Computational Science 2, 718 (2022).
+The authors gratefully acknowledge the LuxProvide teams for their expert support.
+Appendix A Appendix on the Model Architecture
+In analogy with the prediction of the total potential energy as a sum of predicted atomic energies , we predict the total log probability as a sum of per-atom log probabilities (see Equation 2 in the main text).
+In a system with basins, the log probabilities are a vector of length , i.e.
+there is one per-atom logit for each basin.
+We use to denote the GNN features from layer for atom .
+First, we forward the node features into a MLP per layer:
+Next, we concatenate for all GNN layers , and use a final linear layer with weights to compute the per-atom log probabilities :
+Finally, the total log probabilities are obtained by summing the individual
+The phase readout layers are analogous to the existing energy readouts found in conventional GNNs [20, 21, 22] (Figure 1, c), and essentially consist of a multilayer perceptron (MLP) containing at most two nonlinearities and about 1000 weights in total (compared to 100,000 weights in the GNN).
+Appendix B Appendix on Model Training
+The predicted log probabilities are normalized and trained to atomic geometries which are labeled with the correct basin.
+The label is obtained by minimizing the energy of the structure, as explained in Figure 1 in the main manuscript.
+In a classification setting, one often uses the cross-entropy loss function [24].
+We can define the phase label of an atomic geometry as a one-hot vector which has the same length as (i.e.
+the number of basins), and whose components indicate the correct basin of attraction:
+Based on Equation 8, we can write the cross-entropy loss as a simple scalar product:
+The classifier is trained while keeping the GNN and energy readouts fixed, similar to conventional transfer learning approaches.
+Its training data consists of atomic geometries which are labeled with their corresponding basin of attraction (e.g.
+A or B).
+Note that the cross-entropy loss is a function defined on the total log probabilities , not on the specific per-atom log probabilities .
+This implies that the cross-entropy loss will incentivize the model to predict correct total log probabilities, but otherwise introduces no restrictions regarding the specific per-atom log probabilities .
+This makes the model prone to overfitting, especially in the low-data regime.
+To improve this, we considered a number of regularization strategies including label smoothing, an L2 penalty on the weights, and an L1 penalty on logit gradients with respect to the atomic coordinates.
+Their efficacy was highly dependent on both the system under study and their specific weights in the loss function, and we chose not to apply them.
+Instead, we designed a new regularization term which prevents the per-atom log probabilities from becoming strongly positive for any given atomic environment .
+First, note that the global probability for each phase can be computed based on the predicted log probabilities using a softmax function:
+for any class A. We can apply a similar reasoning to the per-atom log probabilities in order to obtain per-atom phase probabilities . Next, consider
+The quantity denotes an expected per-atom logit value.
+In practice, given the exponential dependence between the probabilities and the log probabilities, there will only be one nonzero contribution in the sum in Equation 11, and the expectation value will essentially reduce to the largest (per-atom) logit value.
+Note that the predicted per-atom phase (based on the per-atom log probabilities) can differ from the predicted global phase (based on the total log probabilities).
+Finally, in the loss function, we add an L2 regularization on , which essentially prevents the model from being overconfident in any of the per-atom log probabilities:
+For all experiments in this work, we use .
+Furthermore, note that the actual training of the phase readouts can be made very efficient because the GNN node features can be cached and reused at every step of the optimization since the weights of the pretrained GNN are fixed.
+As a result, training is very efficient, and all SBC models in this work can be trained within one hour on a consumer-level GPU.
+Appendix C Appendix on the Active Learning
+Figure 1 (e) in the main manuscript outlines the active learning workflow in more detail.
+We start by training the phase readouts on equilibrium samples from each of the free energy minima, after which we begin with the active learning loop and perform enhanced sampling.
+We introduce harmonic restraints centered on different values of in order to force the system to transition from one basin to another.
+In each iteration, we position 32 umbrellas along the entire CV range, at three different temperatures as to ensure that both low- and (relevant) high-energy regions are properly sampled.
+In each umbrella, the system is randomly initialized in either of the two minima.
+The samples from the resulting trajectory are labeled according to their basin of attraction, as determined by a simple energy minimization for each sample.
+After this, the model is retrained on all data, and the procedure is repeated.
+See the Supplemental Material (Section S1) for a complete overview of all hyperparameters.
+Appendix D Appendix on the Graph Neural Network
+All calculations in this work employed MACE as base potential [22], although we should emphasize that SBC is independent of the specific GNN architecture [40].
+The interatomic potentials for MIL-53(Al) and alanine dipeptide were trained from scratch using psiflow, at the PBE-D3(BJ) level of theory and with a triple zeta basis set [37].
+DFT evaluations were performed with CP2K 2023.1 for MIL-53(Al) in which the basis set was augmented with plane waves up until a cutoff of 1000 Ry, and ORCA 5.0.4 for alanine dipeptide [41, 42].
+For the proton transfer reaction in the zeolite catalyst, we employed a foundational MACE model [38].
+For the MACE models, we employ node features with and 16 channels each, with a cutoff radius of 6.5 Å.
+All models will be made available upon final publication.
+Appendix E Appendix on the Free Energy Calculations
+Using the log probability difference as CV (see equation 4), we can employ conventional enhanced sampling techniques and evaluate the free energy profile .
+Here, we choose to employ umbrella sampling with harmonic restraints on the CV:
+in which we emphasize that the bias energy contains an implicit dependence on both the atomic coordinates and the unit cell parameters through the learned atomic environment features [23].
+We use the multistate Bennett acceptance ratio (MBAR) equations to combine samples from multiple umbrellas in a statistically optimal manner and obtain free energy profiles as shown in Figure 4 [43, 44, 45, 46].
+Note that while relative stabilities are intrinsic to the system, the actual free energy profiles depend on the particular choice of CV, so the learned and reference free energy profiles do not need to have the same shape (although there exist ways to transform one profile into another [47]).